QUESTION 1

(a) Express 27x^-3 y^2/(3x^-2 y^4)^2 without negative indices.

(b) Solve the following equation for x
2^3x^2 × 2^4x = 2^2/2^-2
1
Expert's answer
2019-08-11T09:55:49-0400

(a) "\frac{27x^{-3}y^2}{(3x^{-2}y^4)^2}=\frac{27x^{-3}y^2}{9x^{-4}y^8}=\frac{3x}{y^6}"


(b) "2^3x^2*2^4x=\frac{2^2}{2^{-2}}"

"2^7x^3=2^4"

"x^3=\frac{1}{2^3}"

"x=\frac{1}{2}"